# DiffSSC: Semantic LiDAR Scan Completion using Denoising Diffusion Probabilistic Models

### v2

@@ Abstract
Perception systems play a crucial role in autonomous driving, incorporating multiple sensors and corresponding computer vision algorithms. 3D LiDAR sensors are widely used to capture sparse point clouds of the vehicle’s surroundings. However, such systems struggle to perceive occluded areas and gaps in the scene due to the sparsity of these point clouds and their lack of semantics. To address these challenges, Semantic Scene Completion (SSC) jointly predicts unobserved geometry and semantics in the scene given raw LiDAR measurements, aiming for a more complete scene representation. Building on promising results of diffusion models in image generation and super-resolution tasks, we propose their extension to SSC by implementing the noising and denoising diffusion processes in the point and semantic spaces individually. To control the generation, we employ semantic LiDAR point clouds as conditional input and design local and global regularization losses to stabilize the denoising process. We evaluate our approach on autonomous driving datasets and our approach outperforms the state-of-the-art for SSC.
@@ I Introduction
To provide dense and semantic scene representations for downstream decision-making and action systems, Semantic Scene Completion (SSC) has been proposed, aimed at jointly predicting missing points and semantics from raw LiDAR point clouds. Given its potential to significantly improve scene representation quality, this task has garnered significant attention in the robotics and computer vision communities. Understanding 3D surroundings is an inherent human ability, developed from observing a vast number of complete scenes in daily life. When humans observe scenes from a single view, they can leverage prior knowledge to estimate geometry and semantics. Drawing inspiration from this capability, the SSC model learns prior knowledge of scenes, $\mathit{P}(\text{scene})$, by estimating the complete scene from partial inputs during training. During inference, new partial inputs captured from the scene serve as the likelihood, $\mathit{P}(\text{observation}|\text{scene})$, and the model finally estimates a reasonable posterior result. Notably, the final estimation is not a unique answer but rather a sample from the posterior distribution, $\mathit{P}(\text{scene}|\text{observation})$. This aligns with human intuition, since humans also infer plausible results from partial inputs, while unobserved parts remain subject to infinite possibilities.
@@
However, most traditional SSC methods are limited to learning the prior distribution of data directly, i.e., training a network to estimate the target output directly from partial inputs. Another approach to learning prior distributions is to estimate residuals. Denoising Diffusion Probabilistic Models (DDPMs) gradually introduce noise into the data in the forward diffusion process and employ a denoiser to learn how to remove these noise residuals. The denoiser iteratively predicts and removes noise, allowing the model to recover high-quality data from pure noise. This mechanism effectively learns the prior distribution of the data, which has the potential to be applied in SSC tasks.
In this work, we propose DiffSSC, a novel SSC approach leveraging DDPMs. As shown in Fig. 1, our method jointly estimates missing geometry and semantics from a scene using raw sparse LiDAR point clouds. During training, the model learns the prior distribution by predicting residuals at different noise intensity levels. These multi-level noisy data are generated from ground truth using data augmentation. In the inference stage, the sparse semantic logits serve as conditional input, and the model generates a dense and semantic scene from pure Gaussian noise through a multi-step Markov process. We model both the point and semantic spaces, designing the forward diffusion and reverse denoising processes to enable the model to learn the scene prior to the semantic point cloud representation. In summary, our key contributions are:
@@
  • We separately model the point and semantic spaces to adapt to the diffusion process.
@@
  • We design local and global regularization losses to stabilize the learning process.
@@ II-A LiDAR Perception
LiDAR is widely used in various autonomous agents for collecting 3D point clouds from the environment. In the past, extensive research was dedicated to employing LiDAR for odometry [1] and mapping [2, 3]. Given the inherent challenges of LiDAR, including data sparsity, noise, and outliers, researchers concentrated on developing filtering algorithms [4] and robust point cloud registration [5] to achieve accurate and efficient LiDAR-SLAM systems. With the advent of deep learning, LiDAR data began to be leveraged for object detection [6] and semantic segmentation [7]. Additionally, unlike dense representations such as images, the sparse nature of LiDAR point clouds presents unique challenges for models. To address these challenges, some researchers focus on estimating the gaps between sweeps and occluded regions from sparse point clouds. This has led to the development of semantic scene completion, an emerging technique in LiDAR perception.
@@ II-B Semantic Scene Completion (SSC)
The task of completion has a long research history. Early efforts in this field focused on filling small holes in shapes to enhance model quality, typically employing continuous energy minimization techniques [8]. With the advent of deep learning, approaches evolved to enable networks to learn extensive geometric shape properties [9], allowing to estimate of entire models from partial inputs. In contrast to shape completion, semantic scene completion (SSC) [10] presents a significantly more complex challenge. Scenes exhibit more intricate geometric structures and encompass a wider range of semantic categories. SSCNet [11] represents the pioneering work that formally defined this task. Since its introduction, various input data modalities, such as occupancy grids [12], images [13], and LiDAR-camera fusion [14], have been explored. Additionally, a wide array of methodologies, including point-voxel mapping [15], transformers [16], bird’s-eye view (BEV) assistance [17], and knowledge distillation [18], have been employed to advance the state of the art in this domain. However, these approaches generally operate on voxelized grids, which poses specific challenges for LiDAR point clouds, as voxelization can introduce quantization errors, leading to a loss of resolution and increased memory usage. In this work, we operate directly on point clouds, offering a more efficient method for handling LiDAR data.
@@ II-C Denoising Diffusion Probabilistic Models
Although diffusion models were originally discovered and proposed in the field of physics, DDPMs [19] was the first to apply this method to generative models. In subsequent research, Rombach et al. [20] introduced latent diffusion models, where the diffusion process is performed in the latent space of the image. This significantly improved computational efficiency and reduced resource consumption, enabling the generation of high-quality and high-resolution images, marking a breakthrough in the field of artistic creation. Beyond artistic applications, several works [21, 22] have also adapted diffusion models for LiDAR perception. These approaches typically project 3D data onto image-based representations, such as range images, allowing methods developed for image domains to be directly applied. Notably, due to the higher demands for accuracy in robotics, controlling the generative process to achieve realistic results remains a significant challenge when applying diffusion models in this field. The recent LiDiff [23] directly applies diffusion models to 3D point clouds for scene completion. However, it still lacks the capability to model and process semantics simultaneously. In this work, we apply DDPM to semantic scene completion, to generate dense and accurate semantic scenes.
@@ III Methodology
Figure 2: The overall pipeline of DiffSSC. The raw LiDAR point cloud is semantically segmented using Cylinder3D [7] to generate initial semantic logits. The semantic point cloud is then upsampled, increasing point density for the diffusion process. These duplicated points undergo forward diffusion and reverse denoising, refining their positions and semantic labels. The original semantic point cloud serves as a conditional input, guiding the generation of points in gaps and occluded areas. To further enhance the generated scene, we introduce a refinement model based on MinkUNet [24, 25, 26, 27], which increases the density of the point cloud.
Given a raw LiDAR point cloud, our objective is to estimate a more complete semantic point cloud, including unobserved points with associated semantic labels within gaps and occluded regions. As illustrated in the Fig. 2, we build a diffusion model based on LiDiff [23], supported by a semantic segmentation module and a refinement module. First, the raw LiDAR point cloud is semantically segmented using a Cylinder3D [7] to generate initial semantic logits. Next, we upsample the semantic point cloud to increase point density for the diffusion process. The duplicated semantic points undergo a forward diffusion and a reverse denoising process to adjust their positions and semantics. Notably, the semantic point cloud also serves as a conditional input for the diffusion model, guiding the generation process. The generated scene includes semantic points located in gaps and occluded areas. To further enhance the quality of the generated scene, we designed a refinement model based on MinkUnet [24, 25, 26, 27] to densify the point cloud.
@@ III-A Denoising Diffusion Probabilistic Models (DDPMs)
Ho et al. [19] introduced DDPMs to produce high-quality images through iterative denoising from Gaussian noise. This promising capability is driven by a residual learning mechanism that efficiently captures the data distribution. Specifically, the process begins with a forward diffusion step, during which noise is gradually added to the target data over $T$ steps. The model is then trained to predict the noise added at each step. By predicting and removing noise at time step $t$, the model generates results that closely approximate the raw data distribution.
@@ III-A1 Forward Diffusion Process
Assuming a sample $\bm{x}_{0}\sim q(\bm{x})$ from a target data distribution, the diffusion process gradually adds noise to $\bm{x}_{0}$ over $T$ steps, producing a sequence $\bm{x}_{1},\ldots,\bm{x}_{T}$. When $T$ is large enough, $q(\bm{x}_{T})$ is approximately equal to a normal distribution $\mathcal{N}(\bm{0},\bm{I})$. The intensity of noise added at each step is defined by the noise factors $\beta_{1},\ldots,\beta_{T}$, which significantly influences the performance of the diffusion model. Specifically, at step $t$, Gaussian noise amplified by $\beta_{t}$ is sampled and added to $\bm{x}_{t-1}$. In  [19], the noise parameter $\beta_{t}$ is determined using a linear schedule, starting from an initial value $\beta_{0}$ and linearly increasing over $T$ steps to a final value $\beta_{T}$. Subsequently, several improved noise schedules have been proposed, such as the cosine schedule [28] and the sigmoid schedule [29]. Due to the inefficiency of adding noise step by step, especially during training, where the noise from different steps can be shuffled, one can simplify this process by sampling $\bm{x}_{t}$ from $\bm{x}_{0}$ without computing the intermediate steps $\bm{x}_{1},\ldots,\bm{x}_{t-1}$. To achieve this, Ho et al. [19] define $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, allowing $\bm{x}_{t}$ to be sampled as:
@@
where $\bm{\epsilon}\sim\mathcal{N}(0,I)$. It is important to note that as $T$ is large enough, $q(\bm{x}_{T})$ approaches $\mathcal{N}(\bm{0},\bm{I})$ because $\bar{\alpha}_{T}$ tends to zero.
@@ III-A2 Reverse Denoising Process
The denoising process reverses diffusion and aim to recover the original sample $\bm{x}_{0}$ from Gaussian noise. This is accomplished by a denoiser, which predicts and removes the noise at each step. The reverse diffusion step can be formulated as:
$\bm{x}_{t-1}=\bm{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\bm{\epsilon}_{\theta}(\bm{x}_{t},t)+\frac{1-\bar{\alpha}_{t-1}}{1-\alpha_{t}}\beta_{t}\mathcal{N}(\bm{0},\bm{I})$ (2)
where $\bm{\epsilon}_{\theta}(\bm{x}_{t},t)$ is the noise predicted from $\bm{x}_{t}$ at step $t$. The process of generating the original data can be formulated as a Markov process that repeatedly calls the denoiser until $t=0$. At this point, the model generates a result that approximates $\bm{x}_{0}$. Due to the denoiser effectively learning the high quality of the data distribution $q(\bm{x}_{T})$, the generated samples are of similarly high quality.
@@ III-B Diffusion Semantic Scene Completion
Regarding the principles of DDPMs, we introduce its application in SSC. To focus on the main components, we assume that primary semantic segmentation has been obtained using Cylinder3D. In the context of the diffusion model, the input is a partial semantic point cloud $\mathcal{X}=\{\bm{x}^{1},\dots,\bm{x}^{N}\}$, where each semantic point $\bm{x}^{i}$ is a tuple of a point position and a semantic probability vector $(\bm{p}^{i},\bm{s}^{i})$. Here, $\bm{p}^{i}\in\mathbb{R}^{3}$ represents the 3D coordinates, and $\bm{s}^{i}\in\Delta^{C-1}=\{\bm{s}\in\mathbb{R}^{C}\mid\sum_{j=1}^{C}s^{j}=1,s^{j}\geq 0\}$ lies in the standard $(C-1)$-dimensional simplex, assuming there are $C$ classes in total. The output estimates complete point cloud $\hat{\mathcal{Y}}=\{\hat{\bm{y}}^{1},\dots,\hat{\bm{y}}^{M}\}$. We generate the reference $\mathcal{Y}=\{\bm{y}^{1},\dots,\bm{y}^{M}\}$ by fusing multiple frames with ground-truth semantic labels and then taking the corresponding region as the input scan $\mathcal{X}$. Our goal is to make the estimated $\hat{\mathcal{Y}}$ as close as possible to the ground truth $\mathcal{Y}$.
As mentioned in Sec. I, by learning scene priors, the model gains the ability to estimate a complete scene (posterior) from partial observations (likelihood). The diffusion model efficiently learns the distribution of the ground truth data, acquiring knowledge of the scene prior. To achieve this, we gradually add noise to the ground truth $\mathcal{Y}$, resulting in $\mathcal{Y}_{1},\dots,\mathcal{Y}_{T}$, until $\mathcal{Y}_{T}$ approximates a Gaussian distribution. This form of data augmentation can be simplified using Eq. 1. However, this approach does not directly apply to semantic scene completion. In Eq. 1, the diffusion process is defined as a combined distribution of the sample $\bm{x}_{0}$ and global noise $\bm{\epsilon}\sim\mathcal{N}(0,I)$, with coefficients $\sqrt{\bar{\alpha}_{t}}$ and $\sqrt{1-\bar{\alpha}_{t}}$ controlling the ratio of noise and sample at different time steps. This mechanism works well in shape completion, as the shapes generally approximate a 3D Gaussian distribution. However, the data distribution in large-scale scenes deviates significantly from Gaussianity, particularly due to varying data ranges across different axes. Applying global noise to the entire scene point cloud as a single entity can obscure important details. Therefore, we use the local point diffusion proposed by Nunes et al. [23], which reformulates the diffusion process as a noise offset $\bm{\epsilon}\in\mathbb{R}^{3+C}$ added locally to each point $\bm{y}^{m}\in\mathcal{Y}$, as shown in the following equation:
$\bm{y}_{t}^{m}=\bm{y}^{m}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$ (3)
where $\bm{\epsilon}$ is not an isotropic Gaussian distribution, because of different scaling for 3D positions and semantics.
Although training on ground truth allows the model to generate high-quality results, the process remains inherently uncontrollable. The goal of SSC is to predict a complete scene from partial input, rather than randomly generating scenes. Therefore, during training, we also use the partial semantic point cloud $\mathcal{X}$ as a conditional input, feeding it into the model to guide the point cloud generation. During training, we load a random step $t\in[0,T]$ at each iteration and compute the corresponding $\mathcal{Y}_{t}$ using Eq. 3. The model is trained to estimate the noise at various intensities with the following loss:
@@
$L_{2}(\mathcal{Y}_{t},\mathcal{X},t)=\left\lVert\bm{\epsilon}-\bm{\epsilon}_{\theta}(\mathcal{Y}_{t},\mathcal{X},t)\right\rVert^{2}$ (4)
In traditional loss design, the predicted result is directly compared to the ground truth. However, our approach uses residual learning, where the model’s output is compared to the residual. Therefore, to generate the final scene, the estimated noise must be removed from the noised scene.
During inference, the model begins denoising from Gaussian noise and iteratively predicts and removes the noise in the samples, ultimately generating a dense semantic scene. Since ground truth is not available during inference, Gaussian noise is generated from $\mathcal{X}$. We duplicate the points in $\mathcal{X}$ to match the quantity in the ground truth, ensuring that there are enough points to perform the diffusion process using Eq. 3. Note that the denoiser also takes the partial semantic point cloud $\mathcal{X}$ as a condition to guide the generation process.
III-C Denoiser Design and Regularization

### v3

@@ Abstract
Perception systems play a crucial role in autonomous driving, incorporating multiple sensors and corresponding computer vision algorithms. 3D LiDAR sensors are widely used to capture sparse point clouds of the vehicle’s surroundings. However, such systems struggle to perceive occluded areas and gaps in the scene due to the sparsity of these point clouds and their lack of semantics. To address these challenges, Semantic Scene Completion (SSC) jointly predicts unobserved geometry and semantics in the scene given raw LiDAR measurements, aiming for a more complete scene representation. Building on promising results of diffusion models in image generation and super-resolution tasks, we propose their extension to SSC by implementing the noising and denoising diffusion processes in the point and semantic spaces individually. To control the generation, we employ semantic LiDAR point clouds as conditional input and design local and global regularization losses to stabilize the denoising process. We evaluate our approach on autonomous driving datasets, and it achieves state-of-the-art performance for SSC, surpassing most existing methods.
@@ I Introduction
To provide dense and semantic scene representations for downstream decision-making and action systems, Semantic Scene Completion (SSC) has been proposed, aimed at jointly predicting missing points and semantics from raw LiDAR point clouds. Given its potential to significantly improve scene representation quality, this task has garnered significant attention in the robotics and computer vision communities. Understanding 3D surroundings is an inherent human ability, developed from observing a vast number of complete scenes in daily life. When humans observe a scene from a single view, they can leverage prior knowledge to infer unseen geometry and semantics. Drawing inspiration from this capability, the SSC model learns prior knowledge of scenes, $\mathit{P}(\text{scene})$, by estimating the complete scene from partial inputs during training. During inference, new partial inputs captured from the scene serve as the likelihood, $\mathit{P}(\text{observation}|\text{scene})$, and the model finally estimates a reasonable posterior result. Notably, the final estimation is not a unique answer but rather a sample from the posterior distribution, $\mathit{P}(\text{scene}|\text{observation})$. This aligns with intuition, since humans also infer plausible results from partial inputs, while the unobserved parts can have multiple possible completions.
@@
However, most traditional SSC methods are limited to learning the prior distribution of data directly, i.e., training a model to estimate the target output directly from partial inputs. Another approach to learning prior distributions is to estimate residuals. Denoising Diffusion Probabilistic Models (DDPMs) gradually inject noise into the data in the forward diffusion process and employ a denoiser to learn how to remove these noise residuals. The denoiser iteratively predicts and removes noise, allowing the model to recover high-quality data from pure noise. This mechanism effectively learns the prior distribution of the data, which has the potential to be applied in SSC tasks.
In this work, we propose DiffSSC, a novel SSC approach leveraging DDPMs. As shown in Fig. 1, our method jointly estimates missing geometry and semantics from a scene using raw sparse LiDAR point clouds. During training, the model learns the prior distribution by predicting residuals at different noise intensity levels. These multi-level noisy data are generated from ground truth using data augmentation. In the inference stage, the sparse semantic logits serve as conditional input, and the model generates a dense semantic scene from pure Gaussian noise through a multi-step Markov process. We model both the point and semantic spaces, designing the forward diffusion and reverse denoising processes to enable the model to learn the scene prior to the semantic point cloud representation. In summary, our key contributions are:
@@
  • We jointly model the noise injection process in both the spatial and semantic domains and design corresponding local and global regularization losses to enhance generation quality.
@@ II-A LiDAR Perception
LiDAR is widely used in various autonomous agents for collecting 3D point clouds from the environment. In the past, extensive research was dedicated to employing LiDAR for odometry [1] and mapping [2, 3]. Given the inherent challenges of LiDAR, including data sparsity, noise, and outliers, researchers concentrated on developing filtering algorithms [4] and robust point cloud registration [5] to achieve accurate and efficient LiDAR-SLAM systems. With the advent of deep learning, researchers began focusing on the semantic properties of LiDAR data, with notable applications in object detection [6] and semantic segmentation [7]. Additionally, unlike dense representations such as images, the sparse nature of LiDAR point clouds presents unique challenges for models. To address these challenges, some researchers focus on estimating the gaps between sweeps and occluded regions from sparse point clouds. This has led to the development of semantic scene completion, an emerging technique in LiDAR perception.
@@ II-B Semantic Scene Completion (SSC)
Semantic scene completion (SSC) aims to jointly infer complex geometric structures and diverse semantic categories of a scene from partial observations. Since its introduction, various input data modalities, such as occupancy grids [8], images [9], and LiDAR-camera fusion [10], have been explored. In parallel, a wide array of methodologies, including transformers [11], bird’s-eye view (BEV) assistance [12], and object-centric modeling [13], have been employed to advance the state of the art in this domain. However, these approaches generally operate on voxelized grids, which poses specific challenges for LiDAR point clouds, as voxelization can introduce quantization errors, leading to resolution loss and increased memory usage. In this work, we operate directly on point clouds, offering a more efficient and resolution-preserving method for handling LiDAR data.
@@ II-C Denoising Diffusion Probabilistic Models
Although diffusion models were originally discovered and proposed in the field of physics, DDPMs [14] were the first to apply this method to generative models. In subsequent research, Rombach et al. [15] introduced latent diffusion models, where the diffusion process is performed in the latent space of the image. This significantly improved computational efficiency and reduced resource consumption, enabling the generation of high-quality and high-resolution images, marking a breakthrough in the field of artistic creation. Beyond artistic applications, diffusion models have been extended to spatiotemporal prediction tasks[16] and LiDAR perception [17, 18], where 3D data is often projected onto range images, allowing methods developed for image domains to be directly applied. Notably, due to the higher demands for accuracy in robotics, controlling the generative process to achieve realistic results remains a significant challenge when applying diffusion models in this field. The recent LiDiff [19] directly applies diffusion models to 3D point clouds for scene completion. However, it still lacks the capability to model and process semantics simultaneously. In this work, we apply DDPM to semantic scene completion, to generate dense and accurate semantic scenes.
@@ III Methodology
Figure 2: The overall pipeline of DiffSSC. The raw LiDAR point cloud is semantically segmented using Cylinder3D [20] to generate initial semantic logits. The semantic point cloud is then upsampled. These duplicated points undergo forward diffusion and reverse denoising. The original semantic point cloud serves as a conditional input, guiding the scene generation. To further enhance the generated scene, we introduce a refinement model based on MinkUNet [21, 22, 23, 24], which increases the density of the point cloud.
Given a raw LiDAR point cloud, our objective is to estimate a more complete semantic point cloud, including unobserved points with associated semantic labels within gaps and occluded regions. As illustrated in the Fig. 2, we build a diffusion model supported by a semantic segmentation module and a refinement module. First, the raw LiDAR point cloud is semantically segmented using a Cylinder3D [20] to generate initial semantic logits. Next, we upsample the semantic point cloud to increase point density for the diffusion process. The duplicated semantic points undergo a forward diffusion and a reverse denoising process to adjust their positions and semantics. Notably, the semantic point cloud also serves as a conditional input for the diffusion model, guiding the generation process. The generated scene includes semantic points located in gaps and occluded areas. To further enhance the quality of the generated scene, we designed a refinement model based on MinkUNet [21, 22, 23, 24] to densify the point cloud.
@@ III-A Denoising Diffusion Probabilistic Models (DDPMs)
Ho et al. [14] introduced DDPMs to produce high-quality images through iterative denoising from Gaussian noise. This promising capability is driven by a residual learning mechanism that efficiently captures the data distribution. Specifically, the process begins with a forward diffusion step, during which noise is gradually injected into the target data over $T$ steps. The model is then trained to estimate the noise injected at each step. By predicting and removing noise at time step $t$, the model generates results that closely approximate the raw data distribution.
@@ III-A1 Forward Diffusion Process
Assuming a sample $\bm{x}_{0}\sim q(\bm{x})$ from a target data distribution, the diffusion process gradually adds noise to $\bm{x}_{0}$ over $T$ steps, producing a sequence $\bm{x}_{1},\ldots,\bm{x}_{T}$. When $T$ is large enough, $q(\bm{x}_{T})$ is approximately equal to a normal distribution $\mathcal{N}(\bm{0},\bm{I})$. The intensity of noise added at each step is determined by the noise intensity factors $\beta_{1},\ldots,\beta_{T}$, which significantly influences the performance of the diffusion model. Specifically, at step $t$, Gaussian noise amplified by $\beta_{t}$ is sampled and added to $\bm{x}_{t-1}$. In [14], the noise parameter $\beta_{t}$ is determined using a linear schedule, starting from an initial value $\beta_{0}$ and linearly increasing over $T$ steps to a final value $\beta_{T}$. Subsequently, several improved noise schedules have been proposed, such as the cosine schedule [25] and the sigmoid schedule [26]. Due to the inefficiency of adding noise step by step, especially during batch loading, where the noise from different steps can be shuffled, one can simplify this process by sampling $\bm{x}_{t}$ from $\bm{x}_{0}$ without computing the intermediate steps $\bm{x}_{1},\ldots,\bm{x}_{t-1}$. To achieve this, Ho et al. [14] define $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, allowing $\bm{x}_{t}$ to be sampled as:
@@
where $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$. It is important to note that when $T$ is large enough, $q(\bm{x}_{T})$ approaches $\mathcal{N}(\bm{0},\bm{I})$ because $\bar{\alpha}_{T}$ tends to zero.
@@ III-A2 Reverse Denoising Process
The denoising process reverses diffusion and aims to recover the original sample $\bm{x}_{0}$ from Gaussian noise. This is accomplished by a denoiser, which estimates and removes the noise at each step. The reverse diffusion step can be formulated as:
$\bm{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\bm{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\bm{\epsilon}_{\theta}(\bm{x}_{t},t)\right)+\sigma_{t}\bm{\epsilon},$
$\text{with }\sigma_{t}^{2}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$ (2)
where $\bm{\epsilon}_{\theta}(\bm{x}_{t},t)$ is the noise estimated from $\bm{x}_{t}$ at step $t$. The process of generating the original data can be formulated as a Markov process that repeatedly calls the denoiser until $t=0$. At this point, the model generates a result that approximates $\bm{x}_{0}$. Due to the denoiser effectively learning the high quality of the data distribution $q(\bm{x}_{T})$, the generated samples are of similarly high quality.
@@ III-B Diffusion Semantic Scene Completion
[…]
As mentioned in Sec. I, by learning scene priors, the model gains the ability to estimate a complete scene (posterior) from partial observations (likelihood). The diffusion model efficiently learns the distribution of the ground truth data, acquiring knowledge of the scene prior. To achieve this, we gradually inject noise into the ground truth $\mathcal{Y}$, resulting in $\mathcal{Y}_{1},\dots,\mathcal{Y}_{T}$, until $\mathcal{Y}_{T}$ approximates a Gaussian distribution. However, the noise injection process in Eq. 1 assumes that $\bm{x}_{0}$ is approximately isotropic, which is not suitable for LiDAR-scanned 3D scenes due to significant scale variations across the three spatial dimensions. While normalization can compress the scene into a more isotropic form, it also leads to a significant loss of fine details. To adapt noise injection for LiDAR-scanned 3D scenes, [19] apply local noise offsets at each point. This ensures that the noise intensity remains consistent across all spatial locations. Similarly, semantic categories in autonomous driving follow a long-tailed distribution [27], indicating that their distribution exhibits anisotropy. Inspired by [19], we adopt a local noise offset strategy for semantic noise injection. Specifically, we first scale the one-hot semantic encoding of the ground truth $\mathcal{Y}$ into the logit domain, then add noise offsets independently to each category’s semantic logit, and subsequently restore the probabilistic semantic distribution via softmax. Combining the spatial and semantic domains, we propose a local anisotropic noise injection mechanism.
$\bm{y}_{t}^{m}=\bm{y}^{m}+\sqrt{1-\bar{\alpha}_{t}}\bm{W}\bm{\epsilon},\bm{W}=\begin{bmatrix}\sigma_{p}\bm{I}_{3}&\bm{0}\\ \bm{0}&\sigma_{s}\bm{I}_{C}\end{bmatrix}$ (3)
Here, we employ an anisotropic scaling matrix $\bm{W}$, controlled by the scaling factors $\sigma_{p}$ and $\sigma_{s}$, to modulate the standard Gaussian noise $\bm{\epsilon}$, ensuring that it appropriately adapts to the scale differences in both the spatial and semantic logits domains. The modulated noise is then injected into the local semantic point $\bm{y}^{m}\in\mathbb{R}^{3+C}$. Given a specific time step $t\in[0,T]$, applying Eq. 3 to all points allows for computing the entire noised scene $\mathcal{Y}_{t}$ in a single step, eliminating the need for intermediate computations. This significantly reduces both memory consumption and computational time.
To enable the model to generate a corresponding complete semantic scene based on the current partial input, we encode the partial semantic point cloud $\mathcal{X}$ as a conditional input, which is then fed into the model to guide the point cloud generation process. Thus, the denoiser integrates the noised scene $\mathcal{Y}_{t}$, the conditional input $\mathcal{X}$, and the time step $t$, which indicates the intensity of noise, to estimate the noise $\bm{\epsilon}_{\theta}(\mathcal{Y}_{t},\mathcal{X},t)$.
Based on our residual learning mechanism, we employ the $L_{2}$ loss to regularize the local discrepancy between the estimated noise and the real noise, rather than comparing the generated scene and the target scene.
@@
$L_{2}=\left\lVert\sqrt{1-\bar{\alpha}_{t}}\bm{W}\bm{\epsilon}-\bm{\epsilon}_{\theta}(\mathcal{Y}_{t},\mathcal{X},t)\right\rVert^{2}$ (4)
Therefore, to generate the final scene, we additionally remove the estimated noise from the sample, followed by transforming the semantic logits into a semantic probability vector via the softmax function.
Besides the local loss $L_{2}$ commonly used in DDPM models,  [19] propose a global regularization for the mean and variance of the estimated noise, enforcing it to follow the statistical properties of the injected noise, i.e., a Gaussian distribution. While first- and second-order moments effectively regularize noise in the spatial domain, the estimated noise in the semantic domain tends to exhibit a skewed distribution due to the long-tailed nature of semantic data in autonomous driving. Therefore, we introduce skewness regularization as the third-order constraint to improve the noise distribution in the semantic domain. Thus, the overall loss is formulated as follows:
$L=L_{2}+\lambda_{p}(L_{\text{p,mean}}+L_{\text{p,var}})+\lambda_{s}(L_{\text{s,mean}}+L_{\text{s,var}}+L_{\text{s,skew}})$
$L_{\text{p,mean}}=\bar{\bm{\epsilon}_{p}}^{2},\quad L_{\text{p,var}}=(\hat{\bm{\epsilon}_{p}}-1)^{2}$
$L_{\text{s,mean}}=\bar{\bm{\epsilon}_{s}}^{2},\quad L_{\text{s,var}}=(\hat{\bm{\epsilon}_{s}}-1)^{2},\quad L_{\text{s,skew}}=\tilde{\bm{\epsilon}_{s}}^{2}$ (5)
where $\bm{\epsilon}_{p}$ and $\bm{\epsilon}_{s}$ correspond to the spatial and semantic domains, respectively, with their contributions weighted by $\lambda_{p}$ and $\lambda_{s}$. Meanwhile, $\bar{\bm{\epsilon}_{\theta}}$, $\hat{\bm{\epsilon}_{\theta}}$, and $\tilde{\bm{\epsilon}_{\theta}}$ represent the mean, standard deviation, and skewness of the estimated noise $\bm{\epsilon}_{\theta}$, respectively. Compared to the local term $L_{2}$, the global term is used to regularize the statistical properties of the noise, ensuring that it approximates a Gaussian distribution.
III-C Denoiser Architecture